Draw the structure that fits the following data. Molecular formula: C6H12O 1H NMR spectral data: 13C NMR spectrum:

PLEASE HELP
Answer #1
Concepts and reason

A compound is given with the molecular formula C6H12O{{\rm{C}}_{\rm{6}}}{{\rm{H}}_{{\rm{12}}}}{\rm{O}} , 1HNMR^1{\rm{H - NMR}} , and 13CNMR^{13}{\rm{C - NMR}} having various chemical shift values. The structure of the compound can be deduced by calculating the degree of unsaturation (DU) in the compound and using the given NMR data.

The number of signals obtained in NMR is also depend on the different type of protons present in the compound. Thee splitting pattern is given by (n+1)\left( {{\rm{n}} + {\rm{1}}} \right) rule and intensity of the peak is shown by the pascals triangle.

Fundamentals

• NMR stands for nuclear magnetic resonance which is a useful spectroscopic technique for determining the type of protons and the number of protons in the compounds.

• A set of protons which are present in same chemical environment, that is, present between the same group of atoms are said to be chemically equivalent protons. All chemically equivalent proton generates 1 signal or 1 peak in 1HNMR^1{\rm{H - NMR}} .

• A set of protons which are present in different chemical environment, that is, present between the different group of atoms are said to be chemically non-equivalent protons. All chemically non-equivalent proton generates distinct signal in 1HNMR^1{\rm{H - NMR}} .

• Chemical shift generally depicts the resonance frequency of the nucleus under various electron distribution environment in a compound under magnetic field.

• Splitting of peak: It is also referred as spin-spin splitting. It follows (n+1)\left( {{\rm{n}} + {\rm{1}}} \right) rule, where n is the number of adjacent protons.

As proton NMR shows different chemical shifts for various protons in a compound similarly, 13CNMR^{13}{\rm{C - NMR}} shows different chemical shifts value for chemically non-equivalent carbons.

Chemical shift range for 13CNMR^{13}{\rm{C - NMR}} is higher than that of proton NMR and the range is from 0250ppm0 - 250\,{\rm{ppm}} .

Degree of unsaturation (DU) can be calculated as follows:

DU=nC+1nH+nNnX2{\rm{DU}} = {n_C} + 1 - \frac{{{n_H} + {n_N} - {n_X}}}{2}

Here,

nC=numberofcarbonatomsnH=numberofcarbonatomsnN=numberofcarbonatomsnX=numberofcarbonatoms\begin{array}{l}\\{n_C} = {\rm{number}}\,{\rm{of}}\,{\rm{carbon}}\,{\rm{atoms}}\\\\{n_H} = {\rm{number}}\,{\rm{of}}\,{\rm{carbon}}\,{\rm{atoms}}\\\\{n_N} = {\rm{number}}\,{\rm{of}}\,{\rm{carbon}}\,{\rm{atoms}}\\\\{n_X} = {\rm{number}}\,{\rm{of}}\,{\rm{carbon}}\,{\rm{atoms}}\\\end{array}

The value of DU can be used to arrive at following predictions about the structure of the compound:

DU Value
0
Prediction for structure
Neither unsaturation nor cyclic structure.
Either one ring or one double bond
Two double

Similarly, for higher values such predictions can be made.

Calculate the degree of unsaturation present in the given compound having molecular formula C6H12O{{\rm{C}}_{\rm{6}}}{{\rm{H}}_{{\rm{12}}}}{\rm{O}} as follows:

DU=6+1122=76=1\begin{array}{c}\\{\rm{DU}} = 6 + 1 - \frac{{12}}{2}\\\\ = 7 - 6\\\\ = 1\\\end{array}

DU is calculated as one hence, either a double bond or a ring structure is present in the given compound.

Interpret the chemical shift values given for 13CNMR^{13}{\rm{C - NMR}} . The chemical shift values in ppm from the given NMR are estimated as follows:

205,52,21,11205,\,52,\,21,\,11

The chemical shift is given for four carbons, out of which for three the values 050ppm0 - 50\,{\rm{ppm}} implies the presence of carbon which is bonded to carbon itself, that is, CC{\rm{C - C}} and the value around 205ppm205\,{\rm{ppm}} implies the presence of carbonyl carbon.

Interpret the chemical shift values given for 1HNMR^1{\rm{H - NMR}} as follows:

0.90ppm(6H,triplet)1.53ppm(4H,multiplet)2.34ppm(1H,multiplet)9.72ppm(1H,doublet)\begin{array}{l}\\0.90\,{\rm{ppm}}\left( {6{\rm{H, triplet}}} \right)\\\\1.53\,{\rm{ppm}}\left( {{\rm{4H, multiplet}}} \right)\\\\2.34\,{\rm{ppm}}\left( {{\rm{1H, multiplet}}} \right)\\\\9.72\,{\rm{ppm}}\left( {1{\rm{H, doublet}}} \right)\\\end{array}

The chemical shift 0.90ppm(6H,triplet)0.90\,{\rm{ppm}}\left( {6{\rm{H, triplet}}} \right) indicates the presence of methyl protons, this implies two methyl groups are present in the structure with the carbon containing one two hydrogens.

The chemical shift 1.53ppm(4H,multiplet)1.53\,{\rm{ppm}}\left( {{\rm{4H, multiplet}}} \right) indicates the presence of methylene protons, this implies two methylene groups are present in the structure with the carbons containing more than two chemically different protons.

2.34ppm(1H,multiplet)2.34\,{\rm{ppm}}\left( {{\rm{1H, multiplet}}} \right) indicates the presence of CH{\rm{CH}} proton which is attached with the carbons having more than two chemically different protons.

9.72ppm(1H,doublet)9.72\,{\rm{ppm}}\left( {1{\rm{H, doublet}}} \right) represents the peak of aldehydic proton.

Write the structure of the compound that best suits which the given data as shown below:

-CH3
HC–CH

Ans:

Hence, the correct structure of the compound is as follows:

-CH3
HC–CH